how would the equation:
10(62+4y)/5+3y=84 simplify to (-25)y=-200
?

Answer by stanbon(75887)

how would the equation:
10(62+4y)/5+3y=84 simplify to (-25)y=-200
--------------------
2(62+4y) + 3y = 84
124 + 8y + 3y = 84
11y = -40
y = -40/11
-25y = +1000/11
-----------------------
Ans: It doesn't.
Cheers,
Stan H.